# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ Abstract
This paper gives a nearly tight characterization of the quantum communication complexity of the permutation-invariant Boolean functions. With such a characterization, we show that the quantum and randomized communication complexity of the permutation-invariant Boolean functions are quadratically equivalent (up to a logarithmic factor). Our results extend a recent line of research regarding query complexity [2, 17, 11] to communication complexity, showing symmetry prevents exponential quantum speedups.
Furthermore, we show the Log-rank Conjecture holds for any non-trivial total permutation-invariant Boolean function. Moreover, we establish a relationship between the quantum/classical communication complexity and the approximate rank of permutation-invariant Boolean functions. This implies the correctness of the Log-approximate-rank Conjecture for permutation-invariant Boolean functions in both randomized and quantum settings (up to a logarithmic factor).
@@ 1 Introduction
Exploring quantum advantages is a key problem in the realm of quantum computing. Numerous work focuses on analyzing and characterizing quantum advantages, such as [6, 14, 25, 21, 29, 47]. It has been known that quantum computing demonstrates a potential exponential speedup to solve certain problems than classical computers, such as Simon’s problem [43] and integer factoring problem [42]. However, for some problems, the quantum speedups can be at most polynomial, including the unstructured search problems [26] and collision finding problems [4]. In light of the aforementioned phenomenon, Aaronson [1] proposed such a problem: How much structure is needed for huge quantum speedups?
@@
Regarding the above problem, there exist two major directions to explore the structure needed for quantum speedups in the query model, which is a complexity model commonly used to describe quantum advantages. On the one hand, Aaronson and Ambainis [2] conjectured the acceptance probability of a quantum query algorithm to compute a Boolean function can be approximated by a classical deterministic algorithm with only a polynomial increase in the number of queries, which is still one of most important conjecture in the field of Boolean analysis. On the other hand, Watrous conjectured that the quantum and randomized query complexities are also polynomially equivalent for any permutation-invariant function [2]. Along this direction, Aaronson and Ambainis [2] initiated the study on the quantum speedup of permutation-invariant functions with respect to query complexity. They demonstrated that a function invariant under full symmetry does not exhibit exponential quantum speedups, even if the function is partial, thereby resolving the Watrous conjecture. (Interested readers may refer to [2] for a more detailed introduction.) Furthermore, Chailloux [17] expanded upon their work by providing a tighter bound and removing a technical dependence of output symmetry. Recently, Ben-David, Childs, Gilyén, Kretschmer, Podder and Wang [11] further proved that hypergraph symmetries in the adjacency matrix model allow at most a polynomial separation between randomized and quantum query complexities. All the above results demonstrated that symmetries break exponential quantum speedups in the query model.
@@
While the study of problem structure in the roles of quantum speedups has obtained considerable attention in the query model, it is a natural question to consider whether we can derive similar results in other computation models. The communication complexity model comes to attention as it is also extensively used to demonstrate quantum advantages. Furthermore, while the exponential gap between quantum and classical communication models has been shown in many works [38, 7, 22, 23, 36], there are also some problems in communication models that demonstrate at most polynomial quantum speedups, such as set disjointness problem [39] and (gap) Hamming distance problem [28, 41, 46, 19]. Therefore, it is a meaningful question to consider how much structure is needed for significant quantum speedups in the communication complexity model. More specifically, while symmetry breaks quantum exponential advantages in the query model, does there exist a similar conclusion in the communication complexity model? In this paper, we investigate a variant of the Watrous conjecture concerning the quantum and randomized communication complexities of permutation-invariant functions as follows. Briefly, a permutation-invariant Boolean function is a function that is invariant under permutations of its inputs (see Definition 2.1 for a formal definition).
Conjecture 1.1 (Communication complexity version of the Watrous Conjecture.).
@@
For any permutation-invariant function $f:[m]^{n}\times[m]^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq Q^{*}(f)^{O(1)}$, where $R(f)$ and $Q^{*}(f)$ are the randomized and quantum communication complexities of $f$, respectively.
@@
Furthermore, we study the Log-rank Conjecture proposed by Lovasz and Saks [34], a long-standing open problem in communication complexity. Despite its slow progress on total functions [12, 35, 31], the conjecture has been shown for several subclasses of total permutation-invariant Boolean functions [16] and $\mathsf{XOR}$-symmetric functions [48]. Lee and Shraibman further proposed the Log-approximate-rank Conjecture, stating that the randomized communication and the logarithm of the approximate rank of the input matrix are polynomially equivalent. Surprisingly, this conjecture was later proven false [20], even for its quantum counterpart [5, 44].
@@
Conjecture 1.2 (Log-rank Conjecture for permutation-invariant functions.).
@@
Conjecture 1.3 (Log-Approximate-Rank Conjecture for permutation-invariant functions.).
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq\left(\log\operatorname{\widetilde{rank}}(f)\right)^{O(1)}$, where $\operatorname{\widetilde{rank}}(f)$ is the approximate rank of the input matrix of $f$ (see Definition 2.4 for a formal definition).
Conjecture 1.4 (Quantum Log-Approximate-Rank Conjecture for permutation-invariant functions.).
@@ 1.1 Our Contribution
To study the communication complexity version of the Watrous conjecture, we start with permutation-invariant Boolean functions, which are essential to analyze general permutation-invariant functions. We show that for any permutation-invariant Boolean function, its classical communication complexity has at most a quasi-quadratic blowup comparing to its quantum communication complexity (Theorem 1.5). Thus, we cannot hope for exponential quantum speedups of permutation-invariant Boolean functions. Additionally, Theorem 1.5 gives a nearly tight bound on the quantum communication complexity. Furthermore, we show that every non-trivial permutation-invariant Boolean function satisfies the Log-rank Conjecture in Theorem 1.6. To resolve the (quantum) Log-Approximate-Rank Conjecture, we investigate the relationship between the quantum/classical communication complexities and the approximate rank of any permutation-invariant Boolean function in Theorem 1.7.
Consider a Boolean function $f$. Let $D(f),R(f),Q(f)$ and $Q^{*}(f)$ be the deterministic communication complexity, randomized communication complexity, quantum communication complexity without prior entanglement, and quantum communication complexity of $f$, respectively. Let $\operatorname{\mathrm{rank}}\left(f\right)$ and $\operatorname{\widetilde{rank}}\left(f\right)$ be the rank and approximate rank of $f$. We summarize our results below [^1] In Theorems 1.5 and 1.7, $\widetilde{O}\left(M\left(f\right)\right)=O\left(M(f)\log^{2}n\log\log n\right)$ for any complexity measure $M$..
@@
For any permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 2.1, the followings hold:
$\Omega\left(m\left(f\right)\right)\leq R(f)\leq\widetilde{O}\left(m\left(f\right)^{2}\right)\leq\widetilde{O}\left(Q^{*}(f)^{2}\right)\text{ and }$
$\Omega\left(m\left(f\right)\right)\leq Q^{*}(f)\leq Q(f)\leq\widetilde{O}\left(m\left(f\right)\right),$
where $m(f)$ is a measure defined in Definition 3.1. Hence, $R(f)\leq\widetilde{O}(Q^{*}(f)^{2})$ for any permutation-invariant function $f$.
The complexity measure $m(\cdot)$ is inspired by the work [24], where Gahzi et al. introduced a complexity measure to capture $R(f)$. It is worth noting that their complexity measure is equivalent up to a fourth power of $R(f)$, while our complexity measure $m(\cdot)$ is quadratically related to $R(f)$ and almost tightly characterizes the quantum communication complexity.
@@
For any permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 2.1, we have
$R(f)=\widetilde{O}\left(\log^{2}\operatorname{\widetilde{rank}}\left(f\right)\right)~{}\text{ and}~{}Q(f)=\widetilde{O}\left(\log\operatorname{\widetilde{rank}}\left(f\right)\right).$
@@ 1.2 Proof Techniques
1.2.1 Lower Bound
We outline our approaches to obtain the lower bound on the quantum communication complexity, rank and approximate rank of permutation-invariant functions below:
@@ 1.2.1 Lower Bound
    Quantum communication complexity and approximate rank: In Theorem 1.5, to prove $Q^{*}(f)=\Omega(m(f))$ for any permutation-invariant function $f$, we use the following two-step reduction (see Lemmas 3.4 and 3.2): First, we reduce the lower bound of the quantum communication complexity of the Exact Set-Inclusion Problem ($\mathsf{ESetInc}$, Definition 2.3) to Paturis’s approximate degree of symmetric functions [37] by the pattern matrix method [40], a well-known method for lower bound analysis in quantum communication complexity. Second, we reduce the lower bound of any permutation-invariant function to the lower bound of $\mathsf{ESetInc}$. In Theorem 1.7, we use a similar method to prove the lower bound of approximate rank: $\log\operatorname{\widetilde{rank}}\left(f\right)=\Omega(m(f))$.
@@
    Rank: In Theorem 1.6, we reduce the lower bound of the rank of total permutation-invariant functions to the lower bound of the rank of some representative function instances, such as the set disjointness problem and the equality problem (see Lemma 4.1).
1.2.2 Upper Bound
We use the following methods to show the upper bounds on the communication complexity of permutation-invariant functions in the randomized, quantum, and deterministic models.
@@ 1.2.2 Upper Bound
    Randomized and quantum models: In Theorem 1.5, to prove $R(f)\leq\widetilde{O}\left(m\left(f\right)^{2}\right)$ for any permutation-invariant function $f$, we first propose a randomized protocol to solve the Set-Inclusion problem ($\mathsf{SetInc}$, Definition 2.3) using a well-suited sampling method according to the parameters of $\mathsf{SetInc}$ (see Lemma 3.9). Afterward, we use this protocol as a subroutine to solve any permutation-invariant function based on binary search (see Theorem 3.3). Furthermore, to prove $Q(f)\leq\widetilde{O}\left(m\left(f\right)\right)$, we use the quantum amplitude amplification technique [13, 27] to speed up the above randomized protocol to solve $\mathsf{SetInc}$ (see Lemma 3.10).
@@
    Deterministic model: In Theorem 1.6, to give an upper bound of deterministic communication complexity of total permutation-invariant functions, we propose a deterministic protocol as follows (see Lemma 4.2): Alice and Bob first share their Hamming weight of inputs, and decide who sends the input to the other party according to the definition of function and the Hamming weight of inputs. The party that has all the information of inputs will output the answer. Combining Lemmas 4.1 (described in Section 1.2.1) and 4.2, Theorem 1.6 can be proved.
1.3 Related Work
The need for structure in quantum speedups has been studied in the query model extensively. Beals, Buhrman, Cleve, Mosca and de Wolf [8] demonstrated that there exists at most a polynomial quantum speedup for total Boolean functions in the query model. Moreover, Aaronson and Ambainis [2] established that even partial symmetric functions do not allow super-polynomial quantum speedups. Chailloux [17] further improved this result to a broader class of symmetric functions. Ben-David, Childs, Gilyén, Kretschmer, Podder and Wang [11] later analyzed the quantum advantage for functions that are symmetric under different group actions systematically. Ben-David [10] established a quantum and classical polynomial equivalence for a certain set of functions satisfying a specific symmetric promise. Aaronson and Ben-David [3] proved that if domain $D$ satisfies $D=\text{poly}(n)$, there are at most polynomial quantum speedups for computing an $n$-bit partial Boolean function.
In terms of communication complexity, there are a few results that imply the polynomial equivalence between quantum and classical communication complexity for several instances of permutation-invariant functions. Examples include $\mathsf{AND}$-symmetric functions [39], Hamming distance problem [28, 18], $\mathsf{XOR}$-symmetric functions [48]. While the above results characterized quantum advantage for a certain class of permutation-invariant Boolean functions, our work provides a systemic analysis of all permutation-invariant Boolean functions.
The study of the Log-rank Conjecture and the Log-Approximate-Rank Conjecture has a rich history. Here, we only survey the results about the Log-rank Conjecture and the Log-Approximate-Rank Conjecture about permutation-invariant Boolean functions. Buhuman and de Wolf [16] verified the correctness of the Log-rank Conjecture for $\mathsf{AND}$-symmetric functions. Combining the results of Razborov [39], Sherstov [40] and Suruga [45], it is implied that the Log-Approximate-Rank Conjecture holds for $\mathsf{AND}$-symmetric functions both in the randomized and quantum settings. Moreover, the result of Zhang and Shi [48] implies the Log-rank Conjecture and the (quantum) Log-Approximate-Rank Conjecture hold for $\mathsf{XOR}$-symmetric functions.

### v2

@@ Abstract
This paper gives a nearly tight characterization of the quantum communication complexity of permutation-invariant Boolean functions. With such a characterization, we show that the quantum and randomized communication complexity of permutation-invariant Boolean functions are quadratically equivalent (up to a polylogarithmic factor of the input size). Our results extend a recent line of research regarding query complexity to communication complexity, showing symmetry prevents exponential quantum speedups. Furthermore, we show that the Log-rank Conjecture holds for any non-trivial total permutation-invariant Boolean function. Moreover, we establish a relationship between the quantum/classical communication complexity and the approximate rank of permutation-invariant Boolean functions. This implies the correctness of the Log-approximate-rank Conjecture for permutation-invariant Boolean functions in both randomized and quantum settings (up to a polylogarithmic factor of the input size).
@@ 1 Introduction
Exploring quantum advantages is a key issue in the realm of quantum computing. Numerous work focuses on analyzing and characterizing quantum advantages, such as [5, 13, 24, 20, 28, 47]. It is known that quantum computers can demonstrate a potential exponential speedup compared with classical computers, such as Simon’s problem [43] and integer factoring [42], for which quantum algorithms exploit the internal structures of the problems (e.g. fast Fourier transform). However, for some highly unstructured problems, such as the unstructured search [25] and collision problems [3], quantum speedups are at most polynomial. In light of the aforementioned phenomenon, Aaronson and Ambainis [1] asked: How much structure is needed for huge quantum speedups?
@@
Regarding the above question, there are two major directions to explore the structure needed for quantum speedups in the query model, a complexity model commonly used to describe quantum advantages. On the one hand, Aaronson and Ambainis [1] conjectured the acceptance probability of a quantum query algorithm to compute a Boolean function can be approximated by a classical deterministic algorithm with only a polynomial increase in the number of queries, which is one of most important conjecture in the field of Boolean analysis. On the other hand, Watrous conjectured that quantum and randomized query complexities are polynomially equivalent for any permutation-invariant function [1]. Along this direction, Aaronson and Ambainis [1] initiated the study on the quantum speedups of permutation-invariant functions for query complexity. They demonstrated that any (partial) function that is invariant under full symmetry does not exhibit exponential quantum speedups, thereby resolving the Watrous conjecture. Furthermore, Chailloux [16] expanded upon their work by providing a tighter bound and removing a technical constraint. Recently, Ben-David, Childs, Gilyén, Kretschmer, Podder, and Wang [10] further proved that hypergraph symmetries in the adjacency matrix model allow at most a polynomial separation between randomized and quantum query complexities. All the above results demonstrated that symmetries break exponential quantum speedups in the query model.
@@
The study of the roles of “structure” in quantum speedups has obtained considerable attention in the query model, which leads us to consider whether we can derive similar results in other computation models. In this paper, we study the communication complexity model introduced by Yao [48]. In the model of communication complexity, the inputs are distributed among two separated parties, each party is assumed to be computationally unbounded. The communication complexity studies the minimum number of bits the players need to exchange to achieve a task. In the model of quantum communication complexity [49], the players are allowed to exchange quantum messages. Quantum communication complexity comes to attention as it is also extensively used to demonstrate quantum advantages. Furthermore, while the exponential gap between quantum and classical communication models has been shown in many works [38, 6, 21, 22, 36], there are also some problems in communication models that demonstrate at most polynomial quantum speedups, such as the Set-Disjointness problem [39] and the (gap) Hamming-Distance-Problem [27, 41, 46, 17]. Therefore, it is intriguing to consider how much structure is needed for significant quantum speedups in the communication complexity model. More specifically, would symmetry also break quantum exponential advantages in the communication complexity model? In this paper, we investigate a variant of the Watrous conjecture concerning the quantum and randomized communication complexities of permutation-invariant functions (1.3). Briefly, a permutation-invariant Boolean function is a function that is invariant under permutations of its inputs. We provide the formal definition below.
Definition 1.1 (Permutation-invariant (PI) functions [23]).
Remark 1.2.
For permutation-invariant functions, the function value depends only on the joint type of the input. Specifically, any permutation-invariant function $f$ in Definition 1.1 depends only on $|x|,|y|$ and $|x\land y|$. Here $|\cdot|$ is the Hamming weight of the binary string, i.e., the number of 1’s in the string, and $|x\land y|$ is the number of $i$ such that $x_{i}=y_{i}=1$.
Conjecture 1.3 (Communication complexity version of the Watrous Conjecture).
@@
Fix $m\in\mathbb{Z^{+}}$. For any permutation-invariant function $f:\left\{0,1,...,m\right\}^{n}\times\left\{0,1,...,m\right\}^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq Q(f)^{O(1)}$, where $R(f)$ and $Q(f)$ are the randomized and quantum communication complexities of $f$, respectively.
@@
Furthermore, we study the Log-rank Conjecture proposed by Lovasz and Saks [34], a long-standing open problem in communication complexity. Despite its difficulty on total functions [11, 35, 30], the conjecture has been shown for several subclasses of total permutation-invariant Boolean functions [15] and $\mathrm{XOR}$-symmetric functions [50]. Lee and Shraibman [32] further proposed the Log-approximate-rank Conjecture, stating that the randomized communication and the logarithm of the approximate rank of the input matrix are polynomially equivalent. However, this conjecture was later proven false [19], even for its quantum analogue [4, 44].
@@
Conjecture 1.4 (Log-rank Conjecture for permutation-invariant functions).
@@
Conjecture 1.5 (Log-Approximate-Rank Conjecture for permutation-invariant functions).
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq\left(\log\operatorname{\widetilde{rank}}(f)\right)^{O(1)}$, where $\operatorname{\widetilde{rank}}(f)$ is the approximate rank of the input matrix of $f$ (see Definition 1.7).
Conjecture 1.6 (Quantum Log-Approximate-Rank Conjecture for permutation-invariant functions).
@@
Definition 1.7 (Approximate rank).
For matrix $M\in\left\{-1,1,*\right\}^{m\times n}$ (‘*’ means undefined entry) and $0\leq\epsilon<1$, we say a real matrix $A$ approximates $M$ with error $\epsilon$ if:
  1. 1)
    $|A_{i,j}-M_{i,j}|\leq\epsilon$ for any $i\in[m],j\in[n]$ such that $M_{i,j}\neq*$;
  2. 2)
    $|A_{i,j}|\leq 1$ for all $i\in[m],j\in[n]$.
Let $\mathcal{M}_{\epsilon}$ be the set of all the real matrices that approximate $M$ with error $\epsilon$. The approximate rank of $M$ with error $\epsilon$, denoted by $\operatorname{\widetilde{rank}}_{\epsilon}(M)$, is the smallest rank among all real matrices in $\mathcal{M}_{\epsilon}$. If $\epsilon=2/3$, we abbreviate $\operatorname{\widetilde{rank}}_{\epsilon}(M)$ as $\operatorname{\widetilde{rank}}(M)$. For any (total or partial) Boolean function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\{-1,1,*\}$, let $\operatorname{\widetilde{rank}}(f)\coloneqq\operatorname{\widetilde{rank}}{\left(M_{f}\right)}$, where $M_{f}$ is the input matrix of $f$.
@@ 1.1 Our Contribution
To study the communication complexity version of the Watrous conjecture, we start with permutation-invariant Boolean functions, which is an important step towards fully resolving 1.3. We show that for every permutation-invariant Boolean function, its classical communication complexity has at most a quasi-quadratic blowup compared to its quantum communication complexity (Theorem 1.8). Thus, we cannot hope for significant quantum speedups of permutation-invariant Boolean functions. Additionally, Theorem 1.8 gives a nearly tight bound on the quantum communication complexity up to a polylogarithmic factor of the input size. Furthermore, we show that every non-trivial permutation-invariant Boolean function satisfies the Log-rank Conjecture in Theorem 1.9. To resolve the (quantum) Log-Approximate-Rank Conjecture, we investigate the relationship between the quantum/classical communication complexities and the approximate rank of any permutation-invariant Boolean function in Theorem 1.10.
Consider a Boolean function $f$. Let $D(f),R(f),Q(f)$ be the deterministic communication complexity, randomized communication complexity, and quantum communication complexity of $f$, respectively. Let $\operatorname{\mathrm{rank}}\left(f\right)$ and $\operatorname{\widetilde{rank}}\left(f\right)$ be the rank and approximate rank of $f$. We summarize our results below.
@@
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 1.1, the followings hold:
$\Omega\left(m\left(f\right)\right)\leq R(f)\leq{O\left(m\left(f\right)^{2}\log^{2}n\log\log n+\log n\right)},$
$\Omega\left(m\left(f\right)\right)\leq Q(f)\leq{O\left(m\left(f\right)\log^{2}n\log\log n+\log n\right)},$
where $m(f)$ is a measure defined in Definition 2.6. Hence, $R(f)\leq{O(Q(f)^{2}\log^{2}n\log\log n+\log n)}$ for any permutation-invariant function $f$.
@@
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 1.1, we have
$R(f)$ $={O\left(\log^{2}\operatorname{\widetilde{rank}}\left(f\right)\log^{2}n\log\log n+\log n\right)},$
$Q(f)$ $={O\left(\log\operatorname{\widetilde{rank}}\left(f\right)\log^{2}n\log\log n+\log n\right)}.$
Remark 1.11.
The relations between $R(f),Q(f),\log\operatorname{\widetilde{rank}}(f)$ are tight in Theorems 1.8 and 1.10 (up to a polylogrithmic factor of the input size), since for the Set-Disjointness Problem, we have $R(f)=\Omega(Q(f)^{2})$ and $Q(f)=\Omega\left(\log\operatorname{\widetilde{rank}}(f)\right)$ [33].
1.2 Related Work
The need for structure in quantum speedups has been studied in the query model extensively. Beals, Buhrman, Cleve, Mosca and de Wolf [7] demonstrated that there exists at most a polynomial quantum speedup for total Boolean functions in the query model. Moreover, Aaronson and Ambainis [1] established that partial symmetric functions also do not allow super-polynomial quantum speedups. Chailloux [16] further improved this result to a broader class of symmetric functions. Ben-David, Childs, Gilyén, Kretschmer, Podder and Wang [10] later analyzed the quantum advantage for functions that are symmetric under different group actions systematically. Ben-David [9] established a quantum and classical polynomial equivalence for a certain set of functions satisfying a specific symmetric promise. Aaronson and Ben-David [2] proved that if domain $D$ satisfies $D=\text{poly}(n)$, there are at most polynomial quantum speedups for computing an $n$-bit partial Boolean function.
In terms of communication complexity, there are a few results that imply the polynomial equivalence between quantum and classical communication complexity for several instances of permutation-invariant functions. Examples include $\mathrm{AND}$-symmetric functions [39], Hamming-Distance problem [27, 17], $\mathrm{XOR}$-symmetric functions [50]. While the above results characterized quantum advantage for a certain class of permutation-invariant Boolean functions, our work provides a systemic analysis of all permutation-invariant Boolean functions.
The study of the Log-rank Conjecture and the Log-Approximate-Rank Conjecture has a rich history. Here, we survey the results of the Log-rank Conjecture and the Log-Approximate-Rank Conjecture about permutation-invariant Boolean functions. Buhrman and de Wolf [15] verified the correctness of the Log-rank Conjecture for $\mathrm{AND}$-symmetric functions. Combining the results of Razborov [39], Sherstov [40] and Suruga [45], it is implied that the Log-Approximate-Rank Conjecture holds for $\mathrm{AND}$-symmetric functions both in the randomized and quantum settings. Moreover, Zhang and Shi [50] showed $\mathrm{XOR}$-symmetric functions satisfy the Log-Rank Conjecture. Chattopadhyayand and Mande [18] further proved that the Log-Approximate-Rank Conjecture holds for $\mathrm{XOR}$-symmetric functions for the first time.
@@ 1.3 Proof Techniques
First, we outline our approaches to obtain the lower bound on the quantum communication complexity, rank and approximate rank of permutation-invariant functions below:
@@
    Quantum communication complexity and approximate rank: In Theorem 1.8, we use the following two-step reduction (see Theorem 3.1 and Lemma 3.3): First, we reduce the lower bound of any permutation-invariant function to the lower bound of Exact Set-Inclusion Problem ($\mathrm{ESetInc}$, a specific instance of permutation-invariant functions, Definition 1.13). Second, we reduce the lower bound of the quantum communication complexity of $\mathrm{ESetInc}$ to Paturis’s approximate degree of symmetric functions [37] by the pattern matrix method [40], a well-known method for lower bound analysis in quantum communication complexity. In Theorem 1.10, we use a similar method to prove the lower bound of approximate rank.
@@
    Rank: In Theorem 1.9, we reduce the lower bound of the rank of total permutation-invariant functions to the lower bound of the rank of some representative function instances, such as the Set-Disjointness Problem and the Equality Problem (see Lemma 4.1).
Moreover, we use the following methods to show the upper bounds on the communication complexity of permutation-invariant functions in the randomized, quantum, and deterministic models.
@@
    Randomized and quantum models: In Theorem 1.8, to prove the randomized upper bound for permutation-invariant functions, we first propose a randomized protocol to solve the Set-Inclusion problem ($\mathrm{SetInc}$, Definition 1.13) using a well-suited sampling method according to the parameters of $\mathrm{SetInc}$ (see Lemma 3.8). Afterwards, we use this protocol as a subroutine to solve any permutation-invariant function based on binary search (see Theorem 3.2). Furthermore, to prove quantum upper bound, we use the quantum amplitude amplification technique [12, 26] to speed up the above randomized protocol to solve $\mathrm{SetInc}$ (see Lemma 3.9). It is also worth noting that in the quantum protocols, Alice and Bob do not need to share prior entanglement.
@@
    Deterministic model: In Theorem 1.9, to give an upper bound on the deterministic communication complexity of total permutation-invariant functions, we propose a deterministic protocol as follows (see Lemma 4.4): Alice and Bob first share the Hamming weight of their inputs, and decide who sends the input to the other party. The party that has all the information about inputs will output the answer.
1.4 Comparison with previous work